OpenStudy (anonymous):

(FAN AND MEDAL) help please!!! Graph each function using at least 3 points for each line. All graphs should be placed together on the same coordinate plane, so be sure to label each line

OpenStudy (anonymous):

@ganeshie8 help me please!!!

OpenStudy (anonymous):

@pinkbubbles @Vocaloid ???

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

help please!!!

OpenStudy (anonymous):

@Vocaloid ???

OpenStudy (anonymous):

@Kash_TheSmartGuy

Vocaloid (vocaloid):

what are the functions?

OpenStudy (anonymous):

i dont know :/

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@DanJS

Vocaloid (vocaloid):

is there any other information provided in the problem?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

k

Vocaloid (vocaloid):

please post everything here

OpenStudy (anonymous):

Student Amber Ben Carter Description Amber shared her photo with 3 people. They continued to share it, so the number of shares increases every day as shown by the function. Ben shared his post with 2 friends. Each of those friends shares with 3 more every day, so the number of shares triples every day. Carter shared his post with 10 friends, who each share with only 2 people each day. Social Media Post Shares f(x) = 3(4)x Day Number of Shares 0 2 1 6 2 18 Carter shared his post with 10 friends, who each share with only 2 people each day.

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

are you there? @Vocaloid ???

Vocaloid (vocaloid):

@Nnesha having a bit of trouble making sense of this problem, the information in the table, function, and word problem aren't really matching up

OpenStudy (anonymous):

@Nnesha have the same problem? @Vocaloid ?

Vocaloid (vocaloid):

maybe take a screenshot?

OpenStudy (anonymous):

ok give a sec

Vocaloid (vocaloid):

is there something you need from me?

Vocaloid (vocaloid):

please tag me in a separate question

OpenStudy (anonymous):

you see the table @Vocaloid ???

Vocaloid (vocaloid):

alright, that makes it much more clear now, let's get started

Vocaloid (vocaloid):

we need to write functions for all three people (Amber, Ben, and Carter) the problem already gives us Amber's function f(x) = 3(4)^x, let's remember that for later now, the general form of an exponential function is f(x) = (starting number)(ratio)^x let's find Ben's function. how many shares did he have on day 0?

OpenStudy (anonymous):

ben 2(3)^x carter 10(2)^x

OpenStudy (anonymous):

right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

so whats nexts? or thats the answer?

OpenStudy (phi):

to graph them, they want you to find 3 points. They give you 3 points for Ben You need points for the other 2

OpenStudy (phi):

You answers 1 and 2 now you are doing part 3. graphing

OpenStudy (phi):

*answered

OpenStudy (phi):

I would use x=0 , 1 and 2 remember any number to the 0 power is 1 for amber (at x=0) \[ 3\cdot 4^0= 3 \cdot 1 = 3\] which means (0,3) is the first point for amber. now do x=1 and x=2

OpenStudy (anonymous):

i need to solve?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

solve is not the word. Evaluate which means, if you have a formula and you "put in a number" for x, you get out a new number (the "y" value) For amber, the formula is f(x)= 3*4^x I showed how to find f(0) (which means the y value when x is 0) you replace x with 0, like this f(0)= 3* 4^0 order of operations: do the exponent first.... 4^0 = 1 (special rule about 0, remember?) then 3*1 which is 3 we get f(0)= 3 which means y is 3 and the (x,y) pair is (0,3) now do that with x=1

OpenStudy (phi):

Can you find the y value when x=1 for Amber?

OpenStudy (anonymous):

y=4?

OpenStudy (phi):

Let's see: f(x)= 3*4^x with x=1 we write f(1) = 3* 4^1 is that 4?

OpenStudy (anonymous):

no

OpenStudy (phi):

remember order of operations \[ 3 \cdot 4^1 \] first do exponents. what is \(4^1\) ?

OpenStudy (phi):

the "little 1" as an exponent as in 4^1 gives 4 (try it on a calculator)

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, and f(1)= 3*4 = 12

OpenStudy (phi):

so the points for amber are (0,3), (1,12) now do x=2 can you do that ?

OpenStudy (anonymous):

no, but let me try

OpenStudy (phi):

write each step

OpenStudy (anonymous):

f(x)= 3*2^x with x=2 we write f(1) = 3* 2^1 2^1= 2?

OpenStudy (phi):

First, what equation do we use for amber?? (always good to start out with the correct one)

OpenStudy (anonymous):

so, i did wrong?

OpenStudy (anonymous):

@phi ???

OpenStudy (phi):

you started with the wrong formula

OpenStudy (phi):

please be careful. write down the formula for Amber. then on the next line. *erase the x* and put 2 in its place. what does the line look like after you do that?

OpenStudy (phi):

look at the question. they give you the formula for Amber. Write it down here.

OpenStudy (anonymous):

F(x)=3(4)^x?

OpenStudy (anonymous):

@phi